OpenStudy (lena772):

Write the equation of the line that is perpendicular to the line y = three halvesx + 4 and passes through the point (-6, 3). y = three halvesx - 1 y = three halvesx + 3 y = -two thirdsx - 1 y = -two thirdsx + 3

OpenStudy (anonymous):

your answer is y= -2/3x -1

OpenStudy (anonymous):

would you like the algebra..

OpenStudy (lena772):

yes plz

OpenStudy (anonymous):

y= 3/2x +4 the negative reciprocal is the perpendicular slope, which is -2/3. The line slope intercept form is y-y1=m(x-x1) so you plug in the points into the x, y, including the negative reciprocal slope. y-3=-2/3(x-(-6)) y-3=-2/3(x+6) y-3=-2/3x -4 +3 +3 y= -2/3x-1